As treasurer, Anthony decided to invest $600 collectd in club dues. Part was invested in 3% account for immediate access with the rest in a long term 7 1/4% account. If the interest for 1 year was $26.50, how much was invested at 7 1/4%
O.K. I am stuck again - here is my best effort
0.03x + 0.0725(600-x)= 26.50
.003x + 43.50 - 0.0725= 26.50
0.0425x=17.00
x=400
This is not the answer my book has. Please tell me what I did wrong and what answer you get.
Thank you very much

Found 2 solutions by rapaljer, shree840:
Answer by rapaljer(4671)

SEE MY NOTE BELOW IN ALL CAPS:
As treasurer, Anthony decided to invest $600 collectd in club dues. Part was invested in 3% account for immediate access with the rest in a long term 7 1/4% account. If the interest for 1 year was $26.50, how much was invested at 7 1/4%

YOU ARE ABSOLUTELY CORRECT--SO FAR! YOU JUST DIDN'T ANSWER THE QUESTION! SO WHAT WAS THE QUESTION?? HOW MUCH WAS INVESTED AT 7 1/4% ? YOUR VALUE OF X REPRESENTS THE AMOUNT INVESTED AT 3%. SUBTRACT YOUR ANSWER, WHICH WAS $400 FROM THE TOTAL OF $600, AND YOU GET $200 AT 7 1/4%.
YOU DID IT! NICE JOB!!

0.03x + 0.0725(600-x)= 26.50
.003x + 43.50 - 0.0725x= 26.50
0.0425x=17.00
x=400 at 3%
$600 - x = $600 - 400 = $200 at 7 1/4%
This is not the answer my book has. Please tell me what I did wrong and what answer you get.

R^2 at SCC

Answer by shree840(260)

.03x+.0725(600-x)=26.50
.03x+43.50-.0725x=26.50
( .0300-.0725)x =26.50
-.0425x =26.50-43.50 = -17
x= -17/-.0425 = 400
Investment at 7 1/4 is 600-400 =200